### v1

# ImagenHub: Standardizing the evaluation of conditional image generation models

Abstract
Recently, a myriad of conditional image generation and editing models have been developed to serve different downstream tasks, including text-to-image generation, text-guided image editing, subject-driven image generation, control-guided image generation, etc. However, we observe huge inconsistencies in experimental conditions: datasets, inference, and evaluation metrics – render fair comparisons difficult. This paper proposes ImagenHub, which is a one-stop library to standardize the inference and evaluation of all the conditional image generation models. Firstly, we define seven prominent tasks and curate high-quality evaluation datasets for them. Secondly, we built a unified inference pipeline to ensure fair comparison. Thirdly, we design two human evaluation scores, i.e. Semantic Consistency and Perceptual Quality, along with comprehensive guidelines to evaluate generated images. We train expert raters to evaluate the model outputs based on the proposed metrics. Our human evaluation achieves a high inter-worker agreement of Krippendorff’s alpha on 76% models with a value higher than 0.4. We comprehensively evaluated a total of around 30 models and observed three key takeaways: (1) the existing models’ performance is generally unsatisfying except for Text-guided Image Generation and Subject-driven Image Generation, with 74% models achieving an overall score lower than 0.5. (2) we examined the claims from published papers and found 83% of them hold with a few exceptions. (3) None of the existing automatic metrics has a Spearman’s correlation higher than 0.2 except subject-driven image generation. Moving forward, we will continue our efforts to evaluate newly published models and update our leaderboard to keep track of the progress in conditional image generation.
1 Introduction
Figure 1: The overview of ImagenHub framework, which consists of the newly curated ImagenHub dataset, ImagenHub library, and ImagenHub evaluation platform and standard.
With the recent development of diffusion models, image generation has quickly become one of the most popular research areas in AI. To enable controllability in the image generation process, a myriad of conditional image generation models have been proposed in the past year. Diverse set of conditions have been attempted to steer the diffusion process. One of the most popular task is text-guided image generation (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022), which aims to ground on a text prompt to generate the corresponding image. Besides that, there are also subject-conditioned image generation (Gal et al., 2022; Ruiz et al., 2023), text-guided image editing (Brooks et al., 2023), multi-subject-conditioned image generation (Kumari et al., 2023), style-guided image generation (Sohn et al., 2023), etc. These different tasks aim to serve different types of downstream applications by enabling subject-level, background-level, style-level controls. The field is evolving at an unprecedented pace and lots of improvements have been reported in the published papers. However, one glaring issue we observed is the published work are highly inconsistent in their experimental setups. To summarize, the inconsistencies mainly come from three aspects, namely dataset, inference and evaluation:
  • Dataset: The existing work curated their own evaluation dataset, which makes the comparison of different models totally incomparable.
  • Inference: Some work requires heavy hyper-parameter tuning and prompt engineering to achieve reasonable performance, which makes the model less robust. Due the tuning effort on different models differe significantly, their comaprison could become unfair.
  • Evaluation: The existing work used different human evaluation protocols and guidelines. This inconsistency renders it impossible to compare human evaluation scores across different methods. Moreover, some of the work either employs a single rater or does not report inter-worker agreement. Thus, the reported results might not be comparable across different papers.
Figure 2: The best and the average model performance in each task
These three inconsistencies make it nearly impossible to track the real progress in the field of the conditional image generation. Such an issue could greatly hinder the development of this field. The desiderata is to build a centralized effort to fairly evaluate every model. More importantly, this effort needs to be continuous to keep up with the evolvement of the field. Our paper aims to serve this purpose to standardize the serving and evaluation of all open-source conditional image generation models. More specifically, ImagenHub consists of the modules listed in Figure 1.
ImagenHub Dataset. We surveyed the existing public evaluation sets for all the generation tasks and then picked diverse instances from them to build our ImagenHub dataset. ImagenHub dataset consists of 7 subsets, each with 100-200 instances. This dataset aims at standardizing the evaluation input to ensure fair comparison for different models.
ImagenHub Inference Library. We built a ImagenHub library111it’s similar to Huggingface libraries (Wolf et al., 2019; von Platen et al., 2022) to evaluate all the conditional image generation models. We ported the highly dispersed codebase from the existing works and then standardized them into a unified format. During inference, we fixed the hyper-parameters and the prompt format to prevent per-instance prompt or hyper-parameter tuning, which makes the inference of different models fair and reproducible. The library is designed to be easily extendable.
ImagenHub Evaluator. We explored different human evaluation metrics and iterated over different versions of the rating guidelines to improve the inter-rater agreement. We settled on two three-valued rating metrics ‘Semantic Consistency’ and ‘Perceptual Quality’ to achieve generally high inter-worker agreement measured by Fleiss Kappa (Fleiss & Cohen, 1973) and Krippendorff’s Alpha  (Krippendorff, 2011). We designed a rating standard to achieve several benefits: (1) Our rating guide is an unambiguous checklist table such that the rater can rate the image with ease. (2) The designed rating guideline is unified on every task type. (3) Sustainability. Since each model is rated individually, previous evaluation results can be reused when new models are added.
We demonstrate our evaluation results in Figure 2, where we show the overall score of the best-performing model and the medium-performing model. Based on our evaluation results in section 5, we made some general observations:
  • The existing models’ performance is generally unsatisfying except for Text-guided Image Generation and Subject-driven Image Generation.
  • We found that evaluation results from the published papers from are generally consistent with our evaluation. 83% of the published result ranking is consistent with our ranking.
  • Automatic evaluation metrics do not correlate well with human preference well except subject-driven image generation. The correlation scores are lower than 0.2.
2 The problem of Conditional Image Generation
The goal of conditional image generation is to predict an RGB image $y\in\mathcal{R}^{3\times H\times W}:\mathcal{Y}$, where $H$ and $W$ are the height and width of the image. The prediction of the target image is given a set of input conditions $X=[c_{1},c_{2},\cdots]$, where $X\in\mathcal{X}$, where $C_{i}$ denotes the i-th condition. In the problem, we aim at learning a prediction function $f:\mathcal{X}\rightarrow\mathcal{Y}$ with deep learning models. Here we mainly consider $f$ parameterized with diffusion models. Here we list a set of tasks we consider in Figure 3, where $c_{i}$ can be represented as text prompt, image mask, subject image, source image, background image, control signal, etc.
Task Definition. We formally define the tasks we consider as follows:
  • Text-guided Image Generation: $y=f(p)$, where $p$ is a text prompt describing a scene. The goal is to generate an image consistent with the text description.
  • Mask-guided Image Editing: $y=f(p,I_{\text{mask}},I_{\text{src}})$, where $I_{\text{mask}}$ is a binarized masked image, and $I_{\text{src}}$ is a source image. The goal is to modify given $I_{\text{src}}$ in the masked region according to $p$.
  • Text-guided Image Editing: $y=f(p,I_{\text{src}})$. It’s similar to Mask-guided Image Editing except that there is no mask being provided, the model needs to figure out the region automatically.
  • Subject-driven Image Generation: $y=f(S,p)$, where $S$ is a set of images regarding a specific subject, which normally ranges from 3-5. The goal is to generate an image according to $p$ regarding the subject $S$.
  • Subject-driven Image Editing: $y=f(S,p,I_{\text{src}})$, where $I_{\text{src}}$ is a source image and $S$ is the subject reference. The goal is to replace the subject in $I_{\text{src}}$ with the given subject $S$.
  • Multi-concept Image Composition: $y=f(S_{1},S_{2},p,I_{\text{src}})$, where $S_{1}$ and $S_{2}$ are two sets of concept images. The goal is to compose them together to generate a new image according to the text description $p$.
  • Control-guided Image Generation: $y=f(I_{\text{control}},p)$, where $I_{\text{control}}$ is the control signal like depth map, canny edge, bounding box, etc. The goal is to generate an image following the low-level visual cues.
Figure 3: The visualization of all the conditional image generation tasks. Here we consider tasks with 1-3 conditions, where $\emptyset$ means empty. The special token [V] and [M] are special identifiers.
3 Related Work
Multimodal Conditional Image Synthesis. Recent works in multimodal conditional image synthesis often rely on pre-trained vision-language models such as CLIP  (Radford et al., 2021) and pre-trained large diffusion models like Stable Diffusion  (Rombach et al., 2022). CLIP bridges the gap between textual descriptions and visual content, while Stable Diffusion is a latent diffusion model trained on a massive dataset. CLIP and Stable Diffusion are now widely used as core components in applications such as text-to-image generation, text-guided image editing  (Patashnik et al., 2021; Meng et al., 2021; Couairon et al., 2022; Mokady et al., 2022; Zhang et al., 2023), and mask-guided image editing  (Avrahami et al., 2022; Nichol et al., 2022). Another interesting research direction is image subject personalization with a few or even one image of the subject. Textual Inversion  (Gal et al., 2022) optimizes a text token vector to represent the subject, while DreamBooth  (Ruiz et al., 2023) finetunes the Stable Diffusion model to learn the new subject concept efficiently. These works foster research in potential applications like subject-driven image generation and editing  (Li et al., 2023a; b; Gu et al., 2023), and even multi-concept image compositions  (Kumari et al., 2023). To explore more control conditions, ControlNet  (Zhang & Agrawala, 2023) proposed the usage of zero convolution on the Stable Diffusion model to support additional guided image control. This work brings up the idea of the control-guided image generation task and inspired later work  (Qin et al., 2023) on improving the control versatility.
AI-generated Image Assessment. Evaluating AI-generated images holistically is a complex and open problem  (Salimans et al., 2016). Researchers have proposed various automatic metrics. In the image quality aspect, Inception score  (Salimans et al., 2016), FID (Heusel et al., 2017) are often used. These methods rely on the statistics on an InceptionNet pre-trained on the ImageNet dataset. Despite the fact that they are widely adapted due to the sensitivity to small changes in the image, those metrics are not an ideal solution as they are biased towards the ImageNet dataset and result in inadequate evaluation  (Borji, 2021). Later works like LPIPS  (Zhang et al., 2018) and DreamSim  (Fu et al., 2023) proposed better ways to measure the perceptual similarity. In the semantic consistency aspect, the CLIP score  (Hessel et al., 2021) is often used to measure the vision-language alignment between the generated image and the prompt. Researchers also worked on alternative methods such as BLIP score  (Li et al., 2022) and ImageReward  (Xu et al., 2023). However, in some tasks like subject-driven image generation and editing, the automatic measurement of semantic consistency is still an open problem. One long-established yet effective approach to assessing AI-generated image performance is to rely on human annotators to assess the visual quality  (Denton et al., 2015; Isola et al., 2017; Meng et al., 2021; Chen et al., 2023). The downside is that it entails a reliance on human judgment, which can introduce subjectivity and potentially limit scalability. To mitigate the downsides, the human evaluation design has to be unambiguous and easy to follow.
4 Method
4.1 Human Evaluation Metrics
Our proposed metric can be used in all seven tasks with the same standard. We adopt two major evaluation metrics, namely semantic consistency $SC$ and perceptive quality $PQ$. These two metrics measure the quality of the generated images from two aspects. The semantic consistency measures how well the generated image is aligned with the condition $X=[c_{1},c_{2},\cdots]$. Specifically, we define the semantic consistency score as:
$SC(y,X)=min\{g(y,c_{1}),g(y,c_{2}),\cdots,g(y,c_{k})\}$ (1)
where $g$ is a modularized function to compute the consistency between $y$ and a single condition $c$. We set $g(y,c_{1})\in[0,0.5,1]$, where 0 means inconsistent, 0.5 means partially consistent and 1 means totally consistent. With this formulation, as long as the output is inconsistent with any of the conditions, the evaluation score should become zero. Otherwise, the aggregation function will pick the lowest consistency score from all the conditions. On the other hand, perceptive quality measures the image quality, i.e. whether the image contains artifacts, is blurry, or has an unnatural sense. we set perceptive quality $PQ\in[0,0.5,1]$, where 0 means extremely poor quality, 0.5 means the image has an acceptable quality and 1 means high quality. In these experiments, each model is rated individually. We train human raters to estimate the $g$ function and $PQ$ function with comprehensive guidelines in  Table 1. We derive $O=\sqrt{SC\times PQ}$ as the overall rating of a model. One benefit of using geometric mean as the design choice is that the rating is penalized when one of the aspect scores is too low. We further studied the configuration in section 5.
Condition 1 Condition 2 Condition 3 SC rating
Inconsistent Any Any 0
Any Inconsistent Any 0
Any Any Inconsistent 0
Partially Consistent Any Any 0.5
Any Partially Consistent Any 0.5
Any Any Partially Consistent 0.5
Mostly Consistent Mostly Consistent Mostly Consistent 1.0
Subjects in image Artifacts Unusual sense PQ rating
Unrecognizable Any Any 0
Any Serious Any 0
Recognizable Moderate Any 0.5
Recognizable Any Moderate 0.5
Recognizable Litte/None Little/None 1.0
Table 1: Rating guideline for computing the SC and PQ score. Detail in subsection A.2
Figure 4: Model performance and variance in each subtask.
4.2 Dataset and available models
We present a standardized dataset for each type of task. The information of the datasets is shown in  Table 10. Some models have different standards of inputs in one task. For example, in the text-guided image editing task, DiffEdit is global description-guided while InstructPix2Pix is instruction-guided. We manually created the equivalent meaning prompts for both methods so they can be aligned. All datasets contain a huge variety of test cases to mimic the diversity in real-life situations. We hosted all of our datasets on HuggingFace dataset for easy access and maintenance. Here we demonstrate all the evaluated models in Table 3.
Task Data Source Inference Dataset size
Text-guided Image Generation
DrawBench (Saharia et al., 2022)
DiffusionDB (Wang et al., 2022)
ABC-6K (Feng et al., 2023)
Ours
197
Mask-guided Image Editing
MagicBrush (Zhang et al., 2023)
Ours
179
Text-guided Image Editing
MagicBrush (Zhang et al., 2023)
Ours
179
Subject-driven Image Generation
SuTI (Chen et al., 2023)
Ours
150
Multi-concept Image Composition
CustomDiffsion (Kumari et al., 2023)
Ours
102
Subject-driven Image Editing DreamEditBench (Li et al., 2023b) 154
Control-guided Image Generation HuggingFace community 150
Model #Params Runtime Keywords
Text-to-Image Generation
Dalle-2 (Ramesh et al., 2022) 3.5B 10s unCLIP, two-stage
Stable Diffusion (Rombach et al., 2022) 0.8B 3s Latent Diffusion
DeepFloydIF (deep floyd.ai, 2023) 4.3B 37s Cascaded Pixel Diffusion
OpenJourney (openjourney.ai, 2023) 0.8B 3s SD, Midjourney data
Stable Diffusion XL (stability.ai, 2023) 2.3B 11s Stable Diffusion, X-Large
Mask-guided Image Editing
BlendedDiffusion (Avrahami et al., 2022) 0.5B 57s Noise Blending, DDPM+CLIP
GLIDE (Nichol et al., 2022) 3.5B 19s CLIP, Diffusion
SD-Inpaint (runwayml, 2023) 1.1B 11s SD, Inpainting training
SDXL-Inpaint (stability.ai, 2023) 2.7B 36s SDXL, Inpainting training
Text-guided Image Editing
SDEdit (Meng et al., 2021) 1.3B 13s SDE Prior
Text2Live (Bar-Tal et al., 2022) 3.1M 36s Zero-shot, Edit layer
DiffEdit (Couairon et al., 2022) 1.3B 29s Mask estimation
Cycle Diffusion (Wu & la Torre, 2023) 1.1B 9s DPM-Encoder, Zero-shot
Prompt-to-Prompt (Mokady et al., 2022) 1.1B 2m Cross-Attention
Pix2PixZero (Parmar et al., 2023) 1.1B 21s Cross-Attention, Zero Prompt
InstructPix2Pix (Brooks et al., 2023) 1.1B 11s SD, synthetic P2P data
MagicBrush (Zhang et al., 2023) 1.1B 7s SD, MagicBrush data
Subject-driven Image Generation
Textual Inversion (Gal et al., 2022) 1.1B 15m Word embedding tuning
DreamBooth (Ruiz et al., 2023) 1.1B 10m Finetuning with preservation loss
DreamBooth-Lora (Hu et al., 2021) 1.1B 8m DreamBooth + Low-Rank Adaptation
SuTI (Chen et al., 2023) 2.5B 30s In-context + Apprenticeship learning
BLIP-Diffusion (Li et al., 2023a) 1.1B 8s Pretrained encoder, Zero-shot
Subject-driven Image Editing
DreamEdit (Li et al., 2023b) 1.1B 8m Dreambooth + Region proposal
PhotoSwap (Gu et al., 2023) 1.1B 7m Dreambooth + Cross-Attention
BLIP-Diffusion (Li et al., 2023a) 1.1B 18s Pretrained encoder, Zero-shot
Multi-concept Image Composition
CustomDiffusion (Kumari et al., 2023) 1.1B 19m Cross-attention updating
DreamBooth (Ruiz et al., 2023) 1.1B 11m Finetuning with preservation loss
TextualInversion (Gal et al., 2022) 1.1B 32m Word embedding tuning
Control-guided Image Generation
ControlNet (Zhang & Agrawala, 2023) 1.4B 8s Zero convolution + Frozen model
UniControl (Qin et al., 2023) 1.4B 23s Multi-task pretraining, Zero-shot
Table 3: Overview of all the evaluated models and their parameter size and runtime. The models are listed in chronological order.
5 Experimental Results
Experiment Setup. All the models either used the default setting from the official implementation or the setting suggested in HuggingFace documentation  (von Platen et al., 2022). We disabled negative prompts and any prompt engineering tricks to ensure a fair comparison. We conducted human evaluation by recruiting participants from Prolific and Amazon Mechanical Turk to rate the images, and our own researchers also took part in the image rating process. We assigned 3 raters for each model and computed the SC score, PQ score, and Overall human score. Then we computed the Fleiss kappa  (Fleiss & Cohen, 1973) for each mentioned score. We also computed Krippendorff’s Alpha  (Krippendorff, 2011), which is expected to yield a higher value than Fleiss kappa. This distinction arises from the nature of the rating categories, with Fleiss’ Kappa assuming nominal categories and Krippendorff’s Alpha accommodating ordinal inputs. Both Fleiss’ Kappa and Krippendorff’s Alpha are bounded within the range of [-1, 1], where the value $>0$ indicates an agreement and closer proximity to 1 indicates a higher degree of agreement.
Results. In Figure 4, we present an overview of the model performance across various tasks. Our findings indicate that the performance of the current models is generally underwhelming, with the exception being Text-guided Image Generation and Subject-driven Image Generation, which has models reaching higher than 0.6 on both SC and PQ average. The detailed report on each model’s performance is shown on  Table 4. We noticed that the overall automated metrics’ correlation with SC score and PQ score in each task is below 0.2 except subject-driven image editing task. Metric values are at  Table 7 and  Table 8.
Model LPIPS ↓ CLIP ↑ $SC_{Avg}$ $PQ_{Avg}$ Overall $\text{Fleiss}_{\bar{O}}$ $\text{Kd}_{\bar{O}}$
Text-guided Image Generation
DeepFloydIF N/A 0.2814 0.65$\pm$0.02 0.62$\pm$0.06 0.59$\pm$0.02 0.32 0.51
Stable Diffusion XL N/A 0.2886 0.62$\pm$0.03 0.64$\pm$0.05 0.59$\pm$0.03 0.37 0.61
Dalle-2 N/A 0.2712 0.58$\pm$0.04 0.62$\pm$0.06 0.54$\pm$0.04 0.27 0.40
OpenJourney N/A 0.2814 0.53$\pm$0.02 0.59$\pm$0.05 0.50$\pm$0.02 0.30 0.47
Stable Diffusion 2.1 N/A 0.2899 0.56$\pm$0.02 0.53$\pm$0.05 0.50$\pm$0.03 0.38 0.50
Mask-guided Image Editing
SDXL-Inpainting 0.14 0.2729 0.49$\pm$0.05 0.51$\pm$0.02 0.37$\pm$0.05 0.50 0.72
SD-Inpainting 0.20 0.2676 0.28$\pm$0.04 0.27$\pm$0.10 0.17$\pm$0.07 0.31 0.49
GLIDE 0.21 0.2578 0.20$\pm$0.05 0.48$\pm$0.06 0.16$\pm$0.05 0.33 0.56
BlendedDiffusion 0.34 0.2594 0.12$\pm$0.03 0.11$\pm$0.03 0.05$\pm$0.02 0.36 0.44
Text-guided Image Editing
MagicBrush 0.19 0.2675 0.51$\pm$0.01 0.65$\pm$0.06 0.47$\pm$0.02 0.44 0.67
InstructPix2Pix 0.27 0.2616 0.29$\pm$0.01 0.70$\pm$0.06 0.27$\pm$0.02 0.55 0.74
Prompt-to-prompt 0.36 0.2674 0.17$\pm$0.05 0.55$\pm$0.09 0.15$\pm$0.06 0.36 0.53
CycleDiffusion 0.21 0.2692 0.17$\pm$0.03 0.56$\pm$0.11 0.14$\pm$0.04 0.41 0.63
SDEdit 0.59 0.2872 0.04$\pm$0.03 0.56$\pm$0.12 0.04$\pm$0.03 0.13 0.13
Text2Live 0.06 0.2628 0.02$\pm$0.01 0.82$\pm$0.04 0.02$\pm$0.02 0.10 0.17
DiffEdit 0.16 0.2425 0.02$\pm$0.01 0.23$\pm$0.04 0.01$\pm$0.01 0.24 0.24
Pix2PixZero 0.57 0.2510 0.01$\pm$0.00 0.48$\pm$0.09 0.01$\pm$0.01 0.37 0.37
Subject-driven Image Generation
SuTI 0.77 0.2895 0.64$\pm$0.11 0.68$\pm$0.08 0.58$\pm$0.12 0.20 0.39
DreamBooth 0.77 0.2847 0.51$\pm$0.08 0.93$\pm$0.02 0.55$\pm$0.11 0.37 0.60
BLIP-Diffusion 0.77 0.2729 0.29$\pm$0.04 0.93$\pm$0.04 0.35$\pm$0.06 0.22 0.39
TextualInversion 0.81 0.2680 0.21$\pm$0.04 0.74$\pm$0.08 0.21$\pm$0.05 0.35 0.52
DreamBooth-Lora 0.82 0.2988 0.07$\pm$0.01 0.82$\pm$0.07 0.09$\pm$0.01 0.29 0.37
Subject-driven Image Editing
PhotoSwap 0.34 0.2846 0.34$\pm$0.02 0.65$\pm$0.04 0.36$\pm$0.02 0.35 0.46
DreamEdit 0.22 0.2855 0.31$\pm$0.03 0.61$\pm$0.03 0.32$\pm$0.03 0.33 0.52
BLIP-Diffusion 0.25 0.2901 0.09$\pm$0.03 0.70$\pm$0.02 0.09$\pm$0.03 0.41 0.47
Multi-concept Image Composition
CustomDiffusion 0.79 0.2929 0.26$\pm$0.01 0.86$\pm$0.05 0.29$\pm$0.01 0.73 0.88
DreamBooth 0.78 0.2993 0.11$\pm$0.02 0.78$\pm$0.02 0.13$\pm$0.02 0.61 0.71
TextualInversion 0.80 0.2548 0.04$\pm$0.01 0.74$\pm$0.05 0.05$\pm$0.01 0.62 0.77
Control-guided Image Generation
ControlNet 0.86 0.2555 0.42$\pm$0.05 0.19$\pm$0.04 0.23$\pm$0.04 0.37 0.57
UniControl 0.90 0.2604 0.38$\pm$0.07 0.20$\pm$0.06 0.23$\pm$0.07 0.36 0.58
Table 4: All the evaluated models from seven core tasks. $\text{Fleiss}_{\bar{O}}$ and $\text{Kd}_{\bar{O}}$, denoting Fleiss’ Kappa and Krippendorff’s alpha for the overall average, respectively. We have more automated metric results in Appendix Table 7.
5.1 Discovery and Insights
Text-Guided Image Generation. We observe that all models are able to generate high-quality images. Regarding semantic consistency, all models have a good understanding of the general prompts, while Stable Diffusion XL is better at understanding complex prompts. For example, it exhibits a high degree of accuracy and detail on the prompt “A panda making latte art.” while other models often misunderstood the prompt as “panda latte art”.
Mask-guided Image Editing. We observed the outputs commonly contain obvious artifacts in the masked region boundaries for Stable Diffusion and GLIDE. While Blended Diffusion and Stable Diffusion XL do not suffer from the same issue, they often produce unrecognizable outputs. Stable Diffusion XL obtains the best results but the overall model performance is still far from satisfactory. Another common issue is that the filled regions can hardly harmonize with the background.
Text-guided Image Editing. One key requirement is to edit the image precisely and keep the background untouched. This requirement is indeed challenging because the network has to understand of editing region from the semantic inputs. We discovered that Prompt-to-Prompt, Pix2PixZero, and SDEdit, despite generating high-quality images, often result in completely different backgrounds. We also spotted that in many cases Text2Live will simply return the input as output, this phenomenon also occasionally happened in other models. For paper claims, our evaluation ranking aligns with the findings from CycleDiffsuion, InstructPix2Pix, MagicBrush, Prompt-to-Prompt, and DiffEdit. We found our evaluation ranking does not align with Pix2PixZero, since their paper only tested on word swapping examples, which is not able to generalize to more complex edits.
Subject-driven Image Generation. Our evaluation results largely align with DreamBooth, BLIP-Diffusion, and SuTI findings. Specifically, Textual inversion struggles to maintain target subject features. DreamBooth can imitate subjects based on images but occasionally resorts to copying learned images. DreamBooth-Lora struggles to generate desired subjects but can follow context prompts. BLIP-Diffusion can mimic target subject features but struggles with details. SuTI maintains high consistency with desired subjects and context, with tolerable artifacts in some cases.
Multi-concept Image Composition. Our evaluations validates that CustomDiffusion is consistently better than the other two models. However, while it learns the given multiple subjects’ features better, it could fail to follow the prompts in many cases, especially on actions and positional words. In contrast, DreamBooth learns the correct subjects in some cases, and TextualInversion almost never learn the correct subjects. Interestingly, in some cases where DreamBooth does not learn the correct subjects, it could still follow the prompts correctly.
Subject-driven Image Editing. It is essential to modify the subject from the source to the target without causing excess changes to the background. Human evaluation is also conducted in DreamEdit for the comparison between Photoswap and DreamEdit, but our rankings differ due to varying evaluation criteria. PhotoSwap can adapt to the target subject from the source naturally in most of the cases, but rarely preserves the background well. DreamEdit maintains the context in most of the cases but sometimes leaves observable distortions at the edge of contextualization. BLIP-Diffusion fails the adaptation most of the time, compromising for a more realistic generation.
Control-guided Image Generation. Our evaluation shows there is no significant difference between the two models in both automatic metrics and human evaluation metrics. While UniControl also reported that there is no significant difference in automatic metrics, our human evaluation results do not align. This can be due to the different evaluation standards and aspects. Nevertheless, it has come to our attention that neither of these models demonstrates a high level of robustness. Scratches often appeared on the generated image.
5.2 Ablation Study
Method of overall human score computation. We set the overall score of the model as the geometric mean of SC and PR score (i.e. $O=\sqrt{SC\times PQ}$). But we also explored the weighted sum setting $O=\alpha\times SC+\beta\times PQ$, where both $\alpha$ and $\beta$ are in [0, 1]. We conducted an experiment and found that the weighted sum setting yields a different ranking in the models. Take the text-guided image editing task as an example, Text2Live outperforms CycleDiffusion in the weighted sum setting even though we found that CycleDiffusion performs better in the human examination. We investigated and found that a majority of results in Text2Live simply return the input as output (in that case SC=0 and PQ=1). We tried adjusting the weightings but it still failed to reflect the actual performance of the model. Thus we decided to use the geometric mean setting to penalize the models.
Setting $O=\sqrt{SC\times PQ}$ $O=0.5SC+0.5PQ$ $O=0.7SC+0.3PQ$
$O_{Sum}$ $O_{Avg}$ $O_{Sum}$ $O_{Avg}$ $O_{Sum}$ $O_{Avg}$
MagicBrush 83.51 0.47 103.75 0.58 98.85 0.55
CycleDiffusion 24.89 0.14 65.25 0.36 51.28 0.29
DiffEdit 1.71 0.01 22.08 0.12 14.38 0.08
Text2Live 4.08 0.02 75.25 0.42 46.75 0.26
Table 5: Comparison on overall human score computation setting.
Design choice of human evaluation metric range. When it comes to human evaluation on a massive scale, it’s essential to create a system that’s easy to understand and quick to use. In this investigation, we undertake an exploration into how different settings in this evaluation method can affect the results. Initially, our approach entailed the utilization of a range encompassing [0, 0.5, 1, 2] for both the Semantic Consistency (SC) score and the Perception Quality (PQ) score, where 2 means the image is perfect. However, this configuration yielded suboptimal results in the Fleiss Kappa. Subsequently, an alternative configuration was employed, narrowing the range to [0, 1] for both the SC and PQ scores. This adjustment, while accommodating a binary classification, was observed to yield values that were overly polarized and extreme in nature. To find the right balance between keeping values in a reasonable range and making sure the evaluation method is reliable, our resolution was to define a range of [0, 0.5, 1] while providing explicit and unambiguous guidelines.
Setting categories = [0, 0.5, 1] categories = [0, 0.5, 1, 2] categories = [0, 1]
$O_{Avg}$ $\text{Fleiss}_{\bar{O}}$ $\text{Kd}_{\bar{O}}$ $O_{Avg}$ $\text{Fleiss}_{\bar{O}}$ $\text{Kd}_{\bar{O}}$ $O_{Avg}$ $\text{Fleiss}_{\bar{O}}$ $\text{Kd}_{\bar{O}}$
BLIP-Diffusion 0.13 0.41 0.47 0.09 0.29 0.50 0.05 0.40 0.40
DreamEdit 0.50 0.33 0.52 0.32 0.26 0.55 0.21 0.38 0.38
PhotoSwap 0.62 0.35 0.46 0.36 0.25 0.50 0.29 0.37 0.37
Table 6: Ablation study to understand the impact of the granularity of SC and PQ.
Conclusion
In this paper, we propose ImagenHub as a continuous effort to unify all the efforts in the conditional image generation into a library to ease the access of these models. We standardize the dataset and evaluation of these models to build our ImagenHub Leaderboard. We hope this leaderboard can deliver more reproducible and fair environment for researchers to visualize the progress in these field. A limitation in this work is the reliance on human raters, which is not only expensive and time-consuming. In the future, we plan to work on developing more generic automatic evaluation methods to approximate the human ratings to help people develop better models.
Ethics Statement
Our work aims to benefit the broader research community by providing a standardized framework for evaluating conditional image generation models. For all the benchmarks, we are committed to the ethical use of them. The datasets used in our work are either publicly available or have been collected and curated exercising the utmost respect for privacy and consent. We will maintain a leaderboard to track latest models and encourage open collaboration and discussion in the field. In human evaluation, we followed the minimum hourly wage of \$11. We also ensure that no personal information is collected and no offensive content is presented during human evaluations.
Appendix A Appendix
A.1 More metrics results
Model LPIPS ↓ DINO ↑ CLIP-I ↑ DreamSim ↓ CLIP ↑
Text-guided Image Generation
DeepFloydIF N/A N/A N/A N/A 0.2814
Stable Diffusion XL N/A N/A N/A N/A 0.2886
Dalle-2 N/A N/A N/A N/A 0.2712
OpenJourney N/A N/A N/A N/A 0.2814
Stable Diffusion N/A N/A N/A N/A 0.2899
Mask-guided Image Editing
SDXL-Inpainting 0.14 0.88 0.93 0.16 0.2729
SD-Inpainting 0.20 0.84 0.91 0.21 0.2676
GLIDE 0.21 0.90 0.94 0.14 0.2578
BlendedDiffusion 0.34 0.79 0.86 0.27 0.2594
Text-guided Image Editing
MagicBrush 0.19 0.87 0.91 0.18 0.2675
InstructPix2Pix 0.27 0.79 0.88 0.26 0.2616
Prompt-to-prompt 0.36 0.68 0.80 0.36 0.2674
CycleDiffusion 0.21 0.82 0.89 0.22 0.2692
SDEdit 0.59 0.55 0.74 0.51 0.2872
Text2Live 0.07 0.95 0.96 0.07 0.2628
DiffEdit 0.16 0.82 0.86 0.26 0.2425
Pix2PixZero 0.57 0.53 0.76 0.50 0.2510
Subject-driven Image Generation
SuTI 0.77 0.69 0.81 0.32 0.2895
DreamBooth 0.77 0.61 0.80 0.40 0.2847
BLIP-Diffusion 0.77 0.63 0.82 0.36 0.2729
TextualInversion 0.81 0.37 0.65 0.61 0.2680
DreamBooth-Lora 0.82 0.36 0.70 0.61 0.2988
Subject-driven Image Editing
PhotoSwap 0.34 0.65 0.83 0.37 0.2846
DreamEdit 0.22 0.74 0.87 0.28 0.2855
BLIP-Diffusion 0.25 0.80 0.89 0.23 0.2901
Multi-concept Image Composition
CustomDiffusion 0.79 0.50 0.70 0.55 0.2929
DreamBooth 0.78 0.39 0.64 0.66 0.2993
TextualInversion 0.80 0.44 0.66 0.65 0.2548
Control-guided Image Generation
ControlNet 0.86 0.29 0.67 0.68 0.2555
UniControl 0.90 0.29 0.70 0.69 0.2604
Table 7: All automated metrics for the evaluated models.
PQ ↑ SC ↑
Task corr(LPIPS) corr(DINO) corr(CLIP-I) corr(DreamSim) corr(CLIP)
Text-guided Image Generation N/A N/A N/A N/A -0.0819
Mask-guided Image Editing -0.0502 -0.0421 -0.0308 -0.0294 -0.0257
Text-guided Image Editing 0.0439 0.0449 0.0765 0.0432 0.0437
Subject-driven Image Generation 0.0905 0.1737 0.1204 0.0943 -0.0117
Subject-driven Image Editing 0.2417 0.1690 0.3770 0.3846 -0.0443
Multi-concept Image Composition -0.1018 -0.1657 -0.0965 0.0140 -0.0107
Control-guided Image Generation 0.1864 0.1971 -0.0985 -0.0045 -0.2056
Table 8: Metrics correlation. we inverted the signs for metric that is the lower the better.
A.2 Human Evaluation Guideline
To standardize the conduction of a rigorous human evaluation, we stipulate the criteria for each measurement as follows:
  • Semantic Consistency (SC), score in range [0, 0.5, 1]. It measures the level that the generated image is coherent in terms of the condition provided (i.e. Prompts, Subject Token, etc.).
  • Perceptual Quality (PQ), score in range [0, 0.5, 1]. It measures the level at which the generated image is visually convincing and gives off a natural sense.
Meaning of Semantic Consistency (SC) score:
  • SC=0 : Image not following one or more of the conditions at all (e.g. not following the prompt at all, different background in editing task, wrong subject in subject-driven task, etc.)
  • SC=0.5 : all the conditions are partly following the requirements.
  • SC=1 : The rater agrees that the overall idea is correct.
Meaning of Perceptual Quality (PQ) score:
  • PQ=0 : The rater spotted obvious distortion or artifacts at first glance and those distorts make the objects unrecognizable.
  • PQ=0.5 : The rater found out the image gives off an unnatural sense. Or the rater spotted spotted some minor artifacts and the objects are still recognizable.
  • PQ=1 : The rater agrees that the resulting image looks genuine.
Raters have to strictly adhere to  Table 9 when rating.
Condition 1 Condition 2 (if applicable) Condition 3 (if applicable) SC rating
no following at all Any Any 0
Any no following at all Any 0
Any Any no following at all 0
following some part following some or most part following some or most part 0.5
following some or most part following some part following some or most part 0.5
following some part or more following some or most part following some part 0.5
following most part following most part following most part 1
Objects in image Artifacts Unusual sense PQ rating
Unrecognizable serious Any 0
Recognizable some Any 0.5
Recognizable Any some 0.5
Recognizable none little or None 1
Table 9: Rating guide checklist table.
Artifacts and Unusual sense, respectively, are:
  • Distortion, watermark, scratches, blurred faces, unusual body parts, subjects not harmonized
  • wrong sense of distance (subject too big or too small compared to others), wrong shadow, wrong lighting, etc.
A.3 Rating Examples
There are some examples when evaluating:
Text-to-Image Generation.
Figure 5: Rating examples on Text-to-Image Generation task.
  • OutputA1: [1, 0.5]. SC=1: Prompt perfectly align. PR=0.5: Minor distortion on the cat’s facial features.
  • OutputB1: [0.5, 0.5]. SC=0.5: 3 dogs appeared instead of 1 cat and 2 dogs. PR=0.5: Minor distortion was found on the animal’s face and the watermark.
  • OutputC1: [1, 0.5]. SC=1: The prompt match perfectly with the image. PR=0.5: Minor distortion on the dog’s facial features.
  • OutputA2: [0.5, 1]. SC=0.5: A sign appeared, but failed to spell the word. PR=1: The image look generally real but with some lighting issues.
  • OutputB2: [0.5, 0.5]. SC=0.5: A sign appeared, but failed to spell the word. PR=0.5: The background looks so unnatural.
  • OutputC2: [0.5, 0]. SC=0.5: A sign appeared, but failed to spell the word. PR=0: Heavy distortion on both text and strong artifacts in the background.
Mask-guided Image Editing.
Figure 6: Rating examples on Mask-guided Image Editing task.
  • OutputA1: [1, 0.5]. SC=1: Clearly generate the soccer. PR=0.5: The edit region does not blend well with the context.
  • OutputB1: [1, 1]. SC=1: Successfully add a soccer. PR=1: The soccer is naturally blended with the context.
  • OutputA2: [0, 0]. SC=0: Generated content can not be regarded as horse. PR=0: The middle left part is not natural.
  • OutputB2: [0.5, 0]. SC=0.5: The object is horse-like but not good. PR=0: The whole image is not natural.
Text-guided Image Editing.
Figure 7: Rating examples on Text-guided Image Editing task.
  • OutputA1: [0, 0]. SC=0: The image does not follow the instruction at all. PR=0: Heavy distortion on the cow’s facial features.
  • OutputB1: [1, 1]. SC=1: The hat gives a nice specular reflection. PR=1: It looks real.
  • OutputA2: [0.5, 0.5]. SC=0.5: The hat exists but does not suit well. PR=0.5: The important object look distorted.
  • OutputB2: [0, 0]. SC=0: The background completely changed. PR=0: The whole image look distorted.
Subject-driven Image Generation.
Figure 8: Rating examples on Subject-driven Image Generation task.
  • OutputA1: [1, 1]. SC=1: The output does match the subject and prompt. PR=1: It looks real.
  • OutputB1: [0, 0]. SC=0: The subject dog missing. PR=0: Serious unusual body.
  • OutputA2: [0, 0.5]. SC=0: The output does not match the subject. PR=0.5: Some unusual sense on the backpack.
  • OutputB2: [0, 0.5]. SC=0: The output does not match the prompt. PR=0.5: Some unusual sense on the desk with the grass.
Multi-concept Image Composition.
Figure 9: Rating examples on Multi-concept Image Composition task.
  • OutputA1: [1, 1]. SC=1: Both subjects accurately represent the intended subjects and the prompt actions match. PR=1: In general the image is natural and as real.
  • OutputB1: [0, 1]. SC=0: Subjects are not correct even though prompt action matches. PR=1: The image looks real.
  • OutputA2: [0.5, 0.5]. SC=0.5: Both subjects accurately represent the intended subjects and but the prompt action doesn’t match. PR=0.5: There is a human face on the chair so the important subject chair looks unrealistic, but do not strongly detract from the image’s overall appearance.
  • OutputB2: [0, 0.5]. SC=0: Subject chair missing. PR=0.5: Minor distortion on the cat body and background.
Subject-driven Image Editing.
Figure 10: Rating examples on Subject-driven Image Editing task.
  • OutputA1: [0.5, 0.5]. SC=0.5: Some details of the subject do not match the token. PR=0.5: Some unnatural sense on the hand features.
  • OutputB1: [0.5, 1]. SC=0.5: The subject does match the token, but removing the hand is an unnecessary edit. PR=1: It looks real.
  • OutputA2: [0, 0.5]. SC=0: Subject clock does not match. PR=0.5: Some distortion on the numbers of the clock.
  • OutputB2: [0, 0.5]. SC=0: Subject clock does not match. PR=0.5: Some distortion on the numbers of the clock and the edit region does not blend well with the context.
Control-guided Image Generation.
Figure 11: Rating examples on Control-guided Image Generation task.
  • OutputA1: [1, 0.5]. SC=1: The generated image perfectly described all the required attributes of the user prompt, and even work. PR=0.5: There are some missing details in the background. The details in the far-side are mostly blurred unnaturally.
  • OutputB1: [1, 0.5]. SC=1: The generated image perfectly described all the required attributes of the user prompt. PR=0.5: The details of golden gate bridge are mostly blurred unnaturally.
  • OutputA2: [0, 0]. SC=0: The output does not correspond to the controlled depth image. PR=0: Heavy distortion on the background and the man.
  • OutputB2: [0, 0]. SC=0: The meaning of the text cannot be obtained from the output. PR=0: Heavy distortion.
A.4 Dataset Distribution
Task Data Source Inference Dataset size
Text-guided Image Generation
DrawBench (Saharia et al., 2022)
DiffusionDB (Wang et al., 2022)
ABC-6K (Feng et al., 2023)
Ours
197
Mask-guided Image Editing
MagicBrush (Zhang et al., 2023)
Ours
179
Text-guided Image Editing
MagicBrush (Zhang et al., 2023)
Ours
179
Subject-driven Image Generation
SuTI (Chen et al., 2023)
Ours
150
Multi-concept Image Composition
CustomDiffsion (Kumari et al., 2023)
Ours
102
Subject-driven Image Editing DreamEditBench (Li et al., 2023b) 154
Control-guided Image Generation HuggingFace community 150
Table 10: All the human evaluation datasets from seven core tasks.
(a) Text-to-Image Generation dataset
(b) Mask-guided/Text-guided Image Editing dataset
(c) Subject-driven Image Generation dataset
(d) Multi-concept Image Composition dataset
(e) Subject-driven Image Editing dataset
(f) Control-guided Image Generation dataset
Figure 12: Objects presented in each task’s prompt text.
A.5 More examples visualization
Figure 13: More samples from our Text-guided Image Generation dataset and ImagenHub outputs.
Figure 14: More samples from our Mask-guided Image Editing dataset and ImagenHub outputs.
Figure 15: More samples from our Text-guided Image Editing dataset and ImagenHub outputs.
Figure 16: More samples from our Subject-Driven Image Generation dataset and ImagenHub outputs.
Figure 17: More samples from our Subject-Driven Image Editing dataset and ImagenHub outputs.
Figure 18: More samples from our Multi-Concept Image Composition dataset and ImagenHub outputs.
Figure 19: More samples from our Control-guided Image Generation dataset and ImagenHub outputs.